# Measuring Gravitational Wave Spectrum from Electroweak Phase Transition and Higgs Self-Couplings

### v1

@@ 1 Introduction
The detection and precise measurement of the stochastic gravitational wave background (SGWB) of cosmological origin provides a unique probe of the early Universe (see, e.g., Caldwell:2022qsj ; LiGong:2024qmt for recent reviews). Recent results from pulsar timing arrays (PTAs) NANOGrav:2023gor ; NANOGrav:2023hde ; EPTA:2023sfo ; Athron:2023mer have revealed the first tantalizing evidence of a nanohertz SGWB. Ground-based interferometers, including Advanced LIGO LIGOScientific:2017vtl , Virgo Kumar:2024bfe , and KAGRA KAGRA:2020tym , explore the higher-frequency band from tens to thousands of hertz, and their latest observations have placed strong upper limits in the absence of a detection LIGOScientific:2025bgj . Looking ahead, space-based interferometers such as LISA LISA:2017pwj ; Robson:2018ifk ; LISACosmologyWorkingGroup:2022jok , Taiji Hu:2017mde ; Ruan:2018tsw ; Wu:2018clg , and TianQin TianQin:2015yph ; TianQin:2020hid ; Luo:2020bls will extend the accessible frequency window into the millihertz range. Among the most intriguing cosmological sources of the SGWB are those arising from first-order phase transitions (FOPTs) in the early Universe (see, e.g., Caprini:2015zlo ; Weir:2017wfa ; Mazumdar:2018dfl ; Caprini:2019egz ; Bian:2021ini ; Athron:2023xlk for reviews), which offer a valuable link between cosmology and particle physics by probing symmetry breaking at different energy scales. Of particular interest is the electroweak phase transition (EWPT) associated with the breaking of the electroweak symmetry. In many extensions of the Standard Model, the EWPT can become strongly first order, generating the out-of-equilibrium conditions required for baryogenesis Morrissey:2012db and simultaneously producing a potentially observable GW signal within the reach of future space-based detectors Caprini:2015zlo .
@@
Previous phenomenological studies of the EWPT have mostly relied on simple signal-to-noise ratio (SNR) calculations, combining the predicted SGWB with detector sensitivity curves to assess the detectability of the signal. In practice, however, experimentally extracting a faint cosmological SGWB from the observational data of a space-based detector is particularly challenging, due to the presence of many noise sources and the contamination from astrophysical foregrounds (see, e.g., Boileau:2020rpg ; Boileau:2021sni ).
@@
Space-based GW detectors must contend with various sources of instrumental noise, dominated by laser frequency fluctuations and supplemented by residual acceleration noise, optical-path disturbances, etc LISA:2017pwj . In ground-based experiments, such challenges can be partly mitigated through the use of a detector network, where cross-correlation between independent data streams suppresses uncorrelated noises Allen:1997ad ; Romano:2016dpx . Although a network of space-based detectors has been proposed Ruan:2020smc ; Cai:2023ywp ; Chen:2024fto , most current studies—including the present work—focus on the single-detector case, which requires a different approach for signal extraction. In this context, time-delay interferometry (TDI) Tinto:2014lxa is employed to cancel the dominant laser frequency noise, which results then in three independent data combinations, e.g., the $A$, $E$, and $T$ channels, each with distinct sensitivity to GWs (see, e.g., Smith:2019wny for a recent overview). Among these, the so-called null channel $T$ Adams:2010vc is particularly valuable: at low frequencies, it is almost insensitive to GWs, allowing it to serve as an internal noise monitor that assists in instrument calibration and in distinguishing signal from noise.
@@
In addition, the stochastic signal of cosmological origin is blended with strong astrophysical foregrounds, most notably the unresolved galactic binaries that form a confusion background and the contribution from extragalactic compact binary coalescences Tinto:2004wu ; Smith:2019wny ; Barack:2004wc ; Caprini:2024hue ; Boileau:2020rpg . Disentangling these different components—and thereby establishing a confident detection of the SGWB from a FOPT—requires a dedicated statistical analysis.  [^1] A further challenge arises from deterministic, individually resolvable signals such as those from galactic binaries, massive black hole binaries and extreme mass-ratio inspirals, which must be jointly analyzed together with stochastic components in the so-called global fit framework Cornish:2005qw ; Littenberg:2023xpl ; Rosati:2024lcs . This formidable task remains under active development and will not be considered in this work.
@@
For phenomenological studies, the interplay between GW observations and collider measurements has become a major focus of research Ramsey-Musolf:2019lsf . In particular, the cubic and quartic self-couplings of the Higgs boson—central to understanding the scalar potential and the mechanism of electroweak symmetry breaking but notoriously difficult to measure at colliders DiVita:2017vrr ; Zabinski:2023jhr —serve as key observables that highlight the complementarity between collider and GW approaches Alves:2018jsw .
In this paper, we develop a framework to probe Higgs self-couplings, among others, through space-based detection of phase-transition-induced SGWBs. Starting from a minimal benchmark Lagrangian, we compute the relevant phase transition parameters and the corresponding GW spectra over the parameter space consistent with phenomenological and theoretical constraints. We then construct a likelihood using simulated GW data in the $A$, $E$, and $T$ channels, and perform parameter estimation with both the Fisher information matrix (FIM) and a Bayesian analysis based on Markov Chain Monte Carlo (MCMC) sampling, incorporating realistic astrophysical backgrounds and detector noise for a Taiji-like mission. The measured spectral uncertainties for the injected signal are subsequently mapped back to the parameter space of the particle physics model, enabling further quantitative predictions for the cubic and quartic Higgs self-couplings.
In this work, we demonstrate the above pipeline using the simplest extension of the Standard Model that includes a real scalar singlet—the xSM Barger:2007im ; Profumo:2007wc ; Profumo:2014opa ; Huang:2017jws . This study extends our previous phenomenological analyses Alves:2018jsw by introducing a more realistic statistical treatment tailored to space-based GW experiments.
The remainder of this paper is organized as follows. In Sec. 2, we give a brief introduction to the xSM and its phase transition dynamics. Sec. 3 introduces the GW spectrum from the EWPT and the associated astrophysical foreground and background models. Sec. 4 describes the detector response, noise modeling, and data simulation procedures. In Sec. 5, we present the statistical inference framework, comparing Fisher forecasts with MCMC results. Sec. 6 applies these methods to the simulated data, mapping spectral constraints onto the xSM parameters and then making predictions for the Higgs self-couplings. Finally, Sec. 7 concludes with a discussion of the implications for future GW missions and collider programs.
@@ 2 The Model
In this section, we introduce the particle physics model and its parameter space defined by five independent parameters. Once the model parameters are fixed, the phase transition parameters can be determined, and these are the quantities that directly enter the prediction of the GW spectrum. Throughout this work, we follow the same convention as in our previous study Alves:2018jsw , to which we refer the reader for further details.
@@
The model “xSM” is the simplest extension to the Standard Model and has been extensively studied in the literature due to its simplicity and its ability to accommodate a strongly first-order phase transition Barger:2007im ; Profumo:2007wc ; Profumo:2014opa ; Huang:2017jws . This model is obtained by coupling the SM Higgs doublet with a real scalar gauge singlet, $S=v_{s}+s$, with the following potential
@@
with its vacuum expectation value $v_{\mathrm{EW}}$, and $G^{+},G^{0}$ the Goldstone modes. Through the two minimization conditions around the vacuum ($v_{\mathrm{EW}}$$v_{s}$), $\mu$ and $b_{2}$ are given by:
@@
In turn, $\lambda$, $a_{1}$, $a_{2}$ can be replaced by the mixing angle, $\theta$, the Higgs mass, $m_{h_{1}}\backsimeq 126\ \text{GeV}$, and the mass of the heavier SM-like scalar $m_{h_{2}}$, from the mass matrix diagonalization:
@@
These five parameters define the parameter space of the xSM model, which is constrained firstly by the theoretical requirement of the Higgs potential and experimental measurements of various physical observables. The former includes boundedness of the potential from below, electroweak vacuum stability at zero temperature, perturbativity, perturbative unitarity, etc. The latter includes precision measurements of the $W$ mass, the oblique electroweak corrections, the Higgs signal strengths, searches for a heavier Higgs-like scalar, etc. These have been extensively studied in the literature and our previous study Alves:2018jsw . Our study in this work is based on the phenomenologically viable parameter space, as characterized by a large data set of parameter space points that pass above constraints, and the readers are referred to the previous work for further details of the phenomenological constraints.
@@ 2.2 Phase Transition Parameters
For the GW spectrum generated by phase transitions, it is not the set of five model parameters in Eq. (2.6) that appear directly, but rather a derived set of phase transition parameters Caprini:2024hue :
@@
In the Bayesian statistical analysis presented later, we do not directly infer the above thermodynamic parameters. Instead, our analysis is performed on a more direct spectral model characterized by two parameters: the overall amplitude $\Omega_{0}$ and the peak frequency $f_{\mathrm{p}}$. This choice avoids the issue of parameter degeneracy that would arise if one were to analyze the thermodynamic parameters directly, as will be discussed in subsequent sections. Once constraints on $\Omega_{0}$ and $f_{\mathrm{p}}$ are obtained, we then employ the mapping procedure introduced later to translate these results into constraints on the thermodynamic parameters, and further into the particle-physics model parameters defined in Eq. (2.6). The detailed relations between these parameter sets, which require the calculation of the finite-temperature effective potential and the associated phase transition dynamics, can be found in our previous work Alves:2018jsw and will not be repeated here.
Finally, we note that there remain significant theoretical uncertainties (see, e.g., Croon:2020cgk ; Athron:2022jyi ; Athron:2023rfq ; Lewicki:2024xan ; Zhu:2025pht ) in determining the above phase transition parameters. These arise from several sources: the construction of the finite-temperature effective potential and nucleation rate calculation, including the issue of gauge dependence for perturbative calculation of the effective potential Patel:2011th ; Arunasalam:2021zrs ; Lofgren:2021ogg ; Hirvonen:2021zej ; and the calculation of the parameters $T_{\mathrm{n}}$, $\alpha$, $\beta/H_{\mathrm{n}}$, and $v_{w}$, where different approaches with varying levels of rigor exist Guo:2021qcq . For simplicity, we do not consider these theoretical uncertainties in the following statistical analysis and we refer to Lewicki:2024xan for a study of this kind based on the FIM analysis.
@@ 3.1 Electroweak Phase Transition
In a first-order phase transition, vacuum bubbles are nucleated, then grow, collide, and merge, driving the universe from a metastable to a stable state Guth:1981uk ; Hindmarsh:2020hop . The expansion and collision of these bubbles disturb the surrounding plasma in the form of sound waves and magneto-hydrodynamic (MHD) turbulence Caprini:2015zlo ; Hindmarsh:2013xza . As a result, the SGWB from a first-order phase transition has three main contributions: (i) bubble collisions Kosowsky:1991ua ; Kosowsky:1992rz ; Kosowsky:1992vn ; Huber:2008hg ; Jinno:2016vai ; Jinno:2017fby , (ii) sound waves Hindmarsh:2013xza ; Hindmarsh:2015qta ; Cai:2023guc ; RoperPol:2023dzg , and (iii) MHD turbulence Kosowsky:2001xp ; Caprini:2009yp ; Binetruy:2012ze ; Gogoberidze:2007an ; Niksa:2018ofa ; RoperPol:2019wvy . However, the bubble-collision contribution is usually negligible compared to sound waves for the EWPT, and that of the MHD turbulence is both subdominant and still uncertain. Thus we consider only the dominant contribution from sound waves.
@@ 4.1 Signal Response and Noise Models
Here, the factor $W(f)$ in the results above arises from the inherent complexities of space-based GW detection. The primary source of detector noise stems from the fluctuation of the laser frequency combined with the unequal lengths of the interferometer arms. It should be suppressed by approximately eight orders of magnitude through the use of TDI Tinto:2004wu , a method that generates complex combinations of raw signals to achieve noise cancellation. For simplicity, we assume equal and fixed arm lengths in this analysis. The application of TDI introduces the factor $W=1-e^{-2if/f_{\ast}}$, where $f_{\ast}=\tfrac{c}{2\pi L}$, $c$ denotes the speed of light, and $L$ represents the arm length of the detectors, which is $3\times 10^{9}\ \text{m}$ for Taiji and $2.5\times 10^{9}\ \text{m}$ for LISA. The response functions for the Taiji and LISA detectors are presented in Fig. 1. Notably, at low frequencies, the $T$-channel exhibits significantly lower sensitivity to GW signals compared to the $A$ and $E$ channels, thus serving as the null-channel for GW signals.
@@ 4.2 Data Simulation and Signal Injection
Figure 2: This figure shows the theoretical PSD curve (blue) and the simulated data in channels AA and $T$ (red points). The curve is computed using Eqs. (4.4) and (4.5), while the red points are generated from Eq. (4.16). The analysis focuses on the frequency range $\left[3\times 10^{-5}\,\text{Hz},\,0.5\,\text{Hz}\right]$, with $\Omega_{\text{ast}}=10^{-8}$ and $\varepsilon=2/3$.

### v2

@@ 1 Introduction
The detection and precise measurement of the stochastic gravitational wave background (SGWB) of cosmological origin provide a unique probe of the early Universe (see, e.g., Caldwell:2022qsj; LiGong:2024qmt for recent reviews). Recent results from pulsar timing arrays (PTAs) NANOGrav:2023gor; NANOGrav:2023hde; EPTA:2023sfo; Athron:2023mer have revealed the first tantalizing evidence of a nanohertz SGWB. Ground-based interferometers, including Advanced LIGO LIGOScientific:2017vtl, Virgo Kumar:2024bfe, and KAGRA KAGRA:2020tym, explore the higher-frequency band from tens to thousands of hertz, and their latest observations have placed strong upper limits in the absence of a detection LIGOScientific:2025bgj. Looking ahead, space-based interferometers such as LISA LISA:2017pwj; Robson:2018ifk; LISACosmologyWorkingGroup:2022jok, Taiji Hu:2017mde; Ruan:2018tsw; Wu:2018clg, and TianQin TianQin:2015yph; TianQin:2020hid; Luo:2020bls will extend the accessible frequency window into the millihertz range. Among the most intriguing cosmological sources of the SGWB are those arising from first-order phase transitions (FOPTs) in the early Universe (see, e.g., Caprini:2015zlo; Weir:2017wfa; Mazumdar:2018dfl; Caprini:2019egz; Bian:2021ini; Athron:2023xlk for reviews), which offer a valuable link between cosmology and particle physics by probing symmetry breaking at different energy scales. Of particular interest is the electroweak phase transition (EWPT) associated with the breaking of the electroweak symmetry. In many extensions of the Standard Model, the EWPT can become strongly first order, generating the out-of-equilibrium conditions required for baryogenesis Morrissey:2012db and simultaneously producing a potentially observable GW signal within the reach of future space-based detectors Caprini:2015zlo.
@@
Previous phenomenological studies of the EWPT have mostly relied on simple signal-to-noise ratio (SNR) calculations, combining the predicted SGWB with detector sensitivity curves to assess the detectability of the signal. In practice, however, experimentally extracting a faint cosmological SGWB from the observational data of a space-based detector is particularly challenging due to the presence of many noise sources and the contamination from astrophysical foregrounds (see, e.g., Boileau:2020rpg; Boileau:2021sni).
@@
Space-based GW detectors must contend with various sources of instrumental noise, dominated by laser frequency fluctuations and supplemented by residual acceleration noise, optical-path disturbances, etc LISA:2017pwj. In ground-based experiments, such challenges can be partly mitigated through the use of a detector network, where cross-correlation between independent data streams suppresses uncorrelated noise Allen:1997ad; Romano:2016dpx. Although a network of space-based detectors has been proposed Ruan:2020smc; Cai:2023ywp; Chen:2024fto, most current studies—including the present work—focus on the single-detector case, which requires a different approach for signal extraction. In this context, time-delay interferometry (TDI) Tinto:2014lxa is employed to cancel the dominant laser frequency noise, resulting in three independent data combinations, e.g., the $A$, $E$, and $T$ channels, each with distinct sensitivity to GWs (see, e.g., Smith:2019wny for a recent overview). Among these, the so-called null channel $T$ Adams:2010vc is particularly valuable: at low frequencies, it is almost insensitive to GWs, allowing it to serve as an internal noise monitor that assists in instrument calibration and in distinguishing signal from noise.
@@
In addition, the stochastic signal of cosmological origin is blended with strong astrophysical foregrounds, most notably the unresolved galactic binaries that form a confusion background and the contribution from extragalactic compact binary coalescences Tinto:2004wu; Smith:2019wny; Barack:2004wc; Caprini:2024hue; Boileau:2020rpg. Disentangling these different components—and thereby establishing a confident detection of the SGWB from a FOPT—requires a dedicated statistical analysis.  [^1] A further challenge arises from deterministic, individually resolvable signals such as those from galactic binaries, massive black hole binaries, and extreme mass-ratio inspirals, which must be jointly analyzed along with stochastic components in the so-called global fit framework Cornish:2005qw; Littenberg:2023xpl; Rosati:2024lcs. This formidable task remains under active development and will not be considered in this work.
@@
For phenomenological studies, the interplay between GW observations and collider measurements has become a major focus of research Ramsey-Musolf:2019lsf. In particular, the cubic and quartic self-couplings of the Higgs boson—central to understanding the scalar potential and the mechanism of electroweak symmetry breaking, but notoriously difficult to measure at colliders DiVita:2017vrr; Zabinski:2023jhr—serve as key observables that highlight the complementarity between collider and GW approaches Alves:2018jsw.
In this paper, we develop a framework to probe Higgs self-couplings, among others, through the space-based detection of phase-transition-induced SGWBs. Starting from a minimal benchmark Lagrangian, we compute the relevant phase transition parameters and the corresponding GW spectra over the parameter space consistent with phenomenological and theoretical constraints. We then construct a likelihood using simulated GW data in the $A$, $E$, and $T$ channels and perform parameter estimation with both the Fisher information matrix (FIM) and a Bayesian analysis based on Markov Chain Monte Carlo (MCMC) sampling, incorporating realistic astrophysical backgrounds and detector noise for a Taiji-like mission. The measured spectral uncertainties for the injected signal are subsequently mapped back to the parameter space of the particle physics model, enabling further quantitative predictions for the cubic and quartic Higgs self-couplings.
In this work, we demonstrate the above pipeline using the simplest extension of the Standard Model that includes a real scalar singlet—the xSM Barger:2007im; Profumo:2007wc; Profumo:2014opa; Huang:2017jws; Goncalves:2024vkj. This study extends our previous phenomenological analyses Alves:2018jsw by introducing a more realistic statistical treatment tailored to space-based GW experiments.
The remainder of this paper is organized as follows. In Sec. 2, we provide a brief introduction to the xSM and its phase transition dynamics. Sec. 3 introduces the GW spectrum from the EWPT and the associated astrophysical foreground and background models. Sec. 4 describes the detector response, noise modeling, and data simulation procedures. In Sec. 5, we present the statistical inference framework, comparing Fisher forecasts with MCMC results. Sec. 6 applies these methods to the simulated data, mapping spectral constraints onto the xSM parameters and then making predictions for the Higgs self-couplings. Finally, Sec. 7 concludes with a discussion of the implications for future GW missions and collider programs.
@@ 2 The Model
In this section, we introduce the particle physics model and its parameter space, defined by five independent parameters. Once the model parameters are fixed, the phase transition parameters can be determined; these are the quantities that directly enter the prediction of the GW spectrum. Throughout this work, we follow the same convention as in our previous study Alves:2018jsw, to which we refer the reader for further details.
@@
The model “xSM” is the simplest extension of the Standard Model and has been extensively studied in the literature due to its simplicity and its ability to accommodate a strongly first-order phase transition Barger:2007im; Profumo:2007wc; Profumo:2014opa; Huang:2017jws. This model is obtained by coupling the SM Higgs doublet with a real scalar gauge singlet, $S=v_{s}+s$, with the following potential
@@
with its vacuum expectation value $v_{\mathrm{EW}}$ and $G^{+},G^{0}$, the Goldstone modes. Through the two minimization conditions around the vacuum ($v_{\mathrm{EW}}$$v_{s}$), $\mu$ and $b_{2}$ are given by:
@@
In turn, $\lambda$, $a_{1}$, and $a_{2}$ can be replaced by the mixing angle $\theta$, the Higgs mass $m_{h_{1}}\backsimeq 126\ \text{GeV}$, and the mass of the heavier SM-like scalar $m_{h_{2}}$, from the mass matrix diagonalization:
@@
These five parameters define the parameter space of the xSM model, which is constrained firstly by the theoretical requirement of the Higgs potential and by experimental measurements of various physical observables. The former includes the boundedness of the potential from below, electroweak vacuum stability at zero temperature, perturbativity, perturbative unitarity, etc. The latter includes precision measurements of the $W$ mass, the oblique electroweak corrections, the Higgs signal strengths, searches for a heavier Higgs-like scalar, etc. These have been extensively studied in the literature and our previous study Alves:2018jsw. Our study in this work is based on the phenomenologically viable parameter space, characterized by a large dataset of parameter space points that pass above the constraints. The readers are referred to the previous work for further details on the phenomenological constraints.
@@ 2.2 Phase Transition Parameters
For the GW spectrum generated by phase transitions, it is not the set of five model parameters in Eq. (2.6) that appears directly, but rather a derived set of phase transition parameters Caprini:2024hue:
@@
In the Bayesian statistical analysis presented later, we do not directly infer the above thermodynamic parameters. Instead, our analysis is performed on a more direct spectral model characterized by two parameters: the overall amplitude $\Omega_{0}$ and the peak frequency $f_{\mathrm{p}}$. This choice avoids the issue of parameter degeneracy that would arise if one were to analyze the thermodynamic parameters directly, as will be discussed in subsequent sections. Once constraints on $\Omega_{0}$ and $f_{\mathrm{p}}$ are obtained, we then employ the mapping procedure introduced later to translate these results into constraints on the thermodynamic parameters and further into the particle-physics model parameters defined in Eq. (2.6). The detailed relations between these parameter sets, which require the calculation of the finite-temperature effective potential and the associated phase transition dynamics, can be found in our previous work Alves:2018jsw and will not be repeated here.
Finally, we note that there remain significant theoretical uncertainties (see, e.g., Croon:2020cgk; Athron:2022jyi; Athron:2023rfq; Lewicki:2024xan; Zhu:2025pht) in determining the above phase transition parameters. These arise from several sources: the construction of the finite-temperature effective potential and the nucleation rate calculation, including the issue of gauge dependence for the perturbative calculation of the effective potential Patel:2011th; Arunasalam:2021zrs; Lofgren:2021ogg; Hirvonen:2021zej; and the calculation of the parameters $T_{\mathrm{n}}$, $\alpha$, $\beta/H_{\mathrm{n}}$, and $v_{w}$, where different approaches with varying levels of rigor exist Guo:2021qcq. For simplicity, we do not consider these theoretical uncertainties in the following statistical analysis, and we refer to Lewicki:2024xan for a study of this kind based on the FIM analysis.
@@ 3.1 Electroweak Phase Transition
In a first-order phase transition, vacuum bubbles are nucleated, then grow, collide, and merge, driving the universe from a metastable to a stable state Guth:1981uk; Hindmarsh:2020hop. The expansion and collision of these bubbles disturb the surrounding plasma in the form of sound waves and magneto-hydrodynamic (MHD) turbulence Caprini:2015zlo; Hindmarsh:2013xza. As a result, the SGWB from a first-order phase transition has three main contributions: (i) bubble collisions Kosowsky:1991ua; Kosowsky:1992rz; Kosowsky:1992vn; Huber:2008hg; Jinno:2016vai; Jinno:2017fby, (ii) sound waves Hindmarsh:2013xza; Hindmarsh:2015qta; Cai:2023guc; RoperPol:2023dzg, and (iii) MHD turbulence Kosowsky:2001xp; Caprini:2009yp; Binetruy:2012ze; Gogoberidze:2007an; Niksa:2018ofa; RoperPol:2019wvy. However, the bubble-collision contribution is usually negligible compared to sound waves for the EWPT, and that of the MHD turbulence is both subdominant and still uncertain. Thus, we consider only the dominant contribution from sound waves.
@@ 4.1 Signal Response and Noise Models
Here, the factor $W(f)$ in the results above arises from the inherent complexities of space-based GW detection. The primary source of detector noise stems from the fluctuation of the laser frequency combined with the unequal lengths of the interferometer arms. It should be suppressed by approximately eight orders of magnitude through the use of TDI Tinto:2004wu, a method that generates complex combinations of raw signals to achieve noise cancelation. For simplicity, we assume equal and fixed arm lengths in this analysis. The application of TDI introduces the factor $W=1-e^{-2if/f_{\ast}}$, where $f_{\ast}=\tfrac{c}{2\pi L}$, $c$ denotes the speed of light, and $L$ represents the arm length of the detectors, which is $3\times 10^{9}\ \text{m}$ for Taiji and $2.5\times 10^{9}\ \text{m}$ for LISA. The response functions for the Taiji and LISA detectors are presented in Fig. 1. Notably, at low frequencies, the $T$-channel exhibits significantly lower sensitivity to GW signals compared to the $A$ and $E$ channels, thus serving as the null-channel for GW signals.
@@ 4.2 Data Simulation and Signal Injection
Figure 2: This figure shows the theoretical PSD curve (blue) and the simulated data in channels $A$ and $T$ (red points). The curve is computed using Eqs. (4.4) and (4.5), while the red points are generated from Eq. (4.16). The analysis focuses on the frequency range $\left[3\times 10^{-5}\,\text{Hz},\,0.5\,\text{Hz}\right]$, with $\Omega_{\text{ast}}=10^{-8}$ and $\varepsilon=2/3$.